OpenStudy (anonymous):

what is the slope of the line whose equation is y-2=3/5(x-1)?

OpenStudy (anonymous):

can anyone help me?

OpenStudy (anonymous):

3/5

OpenStudy (anonymous):

how did you get it?

OpenStudy (anonymous):

I was thinking that, but wasn't sure, so I didn't post. but if so, what's with x-1? based on y = mx + b (to bnb)

OpenStudy (anonymous):

general formula for a line that passing through point (x1 , y1) and have slope of m is >>> m = (y - y1)/(x-x1)

OpenStudy (anonymous):

your equation is (y-2)/(x-1) = 3/5

OpenStudy (anonymous):

so thats the equation of a line passing point (1,2) and have a slope of 3/5